An educational psychologist does an experiment to compare three different methods of teaching math. Subjects are...

  1. An educational psychologist does an experiment to compare three different methods of teaching math. Subjects are randomly assigned to the different treatments, then are instructed by the relevant method for one month. After completing instruction, all three groups take the same test of achievement. A summary of the results are tabled below. Test whether there is a difference among the 3 methods, summarizing your results in an ANOVA table and in English.

Method A

Method B

Method C

Mean =

80

70

67

Variance =

25

30

20

n =

30

30

30

  1. State your hypotheses.
  1. Determine the characteristics of the comparison distribution.

  1. Determine the cut-off value for a 5% significance level.
  2. Compute the observed F ratio. Summarize results in an ANOVA table.
  1. Decide whether or not to reject the null and give a verbal explanation.
  1. Calculate and interpret h2.
  1. Compute the power of the test. How many subjects are needed to detect an effect with a power of .80 and the given effect size?
  1. Compute a planned comparison between Method A and Method C. Given the number of comparisons that you plan to make, you apply the Bonferroni correction, which suggests testing at the .01 level. What are your conclusions?

Assume now that the educational psychologist had no a priori hypotheses about the nature of any differences. Instead, the researcher wants to compute all possible pairwise comparisons post-hoc. Compute the appropriate post-hoc comparison between Methods A and B (assuming that all possible pairwise comparisons are being computed as well). What are your conclusions?
Answer #1

a.~Let~\mu_i=true~mean~score~due~to~Method~i,~i=A,B,C.\\\\ Null~hypothesis,~H_0:\mu_A=\mu_B=\mu_C~vs.~\\ Alternative~hypothesis,~H_1:~At~least~one~\mu_i~is~different~from~others.\\\\ b.~Comparison~distribution~is~F~distribution.~Here~the~degrees~of~freedom~(df)~of~F~distribution\\ is~(3,87)~and~F~distribution~is~positively~skewed~and~critical~region~lies~right~side~of~F-distribution.\\\\ c.~n=total~no.~of~observations=30*3=90\\ k=no.~of~groups=3\\ Cut-off~value=F_{0.05,3,87}=2.7094~(R~code:~"round(qf(1-0.05,3,87),4)").\\\\

d.~SSE=sum~of~squares~due~to~error=29*(25+30+20)=2175~\\with~df=87.\\\\ MSE=mean~sum~of~squares~due~to~error=2175/87=25.\\\\ \bar{x}=grand~mean=(80+70+67)/3=72.3333\\\\ SSB=sum~of~squares~between~methods\\\\ =30*((80-72.3333)^2+(70-72.3333)^2+(67-72.3333)^2)=2780\\ with~df=3-1=2.\\\\ MSB=mean~sum~of~squares~between~methods=2780/2=1390\\\\ F~ratio=1390/25=55.6\\\\ ANOVA~table\\\\ SST=sum~of~squares~due~to~total=SSE+SSB=2175+2780=4955\\\\ \begin{matrix} Source &SS & df& MS &F \\ Between&2780 & 2 &1390 &55.6 \\ Error& 2175 &87 &25 & \\ Total&4955 & 89 & & \end{matrix}\\\\ Since~value~of~F>Cut-off~value,~we~reject~H_0~at~5\%~level~of~significance.\\\\ e.~\eta^2=\frac{SSB}{SST}=2780/4955=0.56~and~large~effect~size.